OpenStudy (anonymous):

Find the number of real solutions for the equation 3x^2_6x+3=0

OpenStudy (anonymous):

3x^2-6x+3=0 ?

OpenStudy (anonymous):

whoops +3

OpenStudy (anonymous):

3x+6x+3=0

OpenStudy (anonymous):

ax^2+bx+c = 0 \[\Delta=b^2-4ac\] if \[\Delta = 0\] then one real root if \[\Delta > 0\] two real roots if \[\Delta < 0\] no real roots

OpenStudy (anonymous):

this is called discriminant

OpenStudy (anonymous):

so there is none?

OpenStudy (anonymous):

or two?

OpenStudy (anonymous):

3x^2+6x+3=0 this is the equation ?

OpenStudy (anonymous):

3x+6x+3=0

OpenStudy (anonymous):

no square ?

OpenStudy (anonymous):

so what i said isnt relevant

OpenStudy (anonymous):

9x + 3 =0 9x = -3 x = -3/9 one solution here

OpenStudy (anonymous):

but are you sure there is no x^2 ?!

OpenStudy (anonymous):

positive thanks.